OpenStudy (redprince):

2x+y-3=0 x+2y+3=0 is these equations are solve like this:

OpenStudy (anonymous):

Multiply first equation by 2 4x + 2y - 6 = 0 x + 2y + 3 = 0 Subtract them: 3x = 9 x = 3 Put the value of x in any of the above equation and find the value of y.. Can you do it or not??

OpenStudy (redprince):

@waterineyes I want to solve it simultaneously: look \[x/(1)(-3) + (2)(3) = -y/(-3)(2) + (1)(3) = 1/(2)(2) +(1)(1)\]

OpenStudy (anonymous):

yes waterineyes do it perfecly

OpenStudy (anonymous):

@KRAZZy What???

OpenStudy (redprince):

@KRAZZy I know and I also solve it! But I forgot that how to solve equations simultaneously!

OpenStudy (anonymous):

I studied this method but now forgot how to solve by this method but I will try wait for sometime...

OpenStudy (redprince):

@waterineyes O.k., I am waiting for your reply!!!

OpenStudy (anonymous):

can any1 tell me wht is simultaneous methd:P

OpenStudy (zepp):

Are you referring to the reduction method?

OpenStudy (zepp):

If so (x+2y+3=0)*2 = 2x+4y+6=0

OpenStudy (zepp):